# A Comprehensive Survey on Machine Learning Driven Material Defect Detection: Challenges, Solutions, and Future Prospects

### v1

@@ 3.1.3. Principal Component Analysis
Principal Component Analysis (PCA) is a statistical technique used in material defect detection to simplify the complexity of high-dimensional data while retaining trends and patterns (cao2024crack, ). Active thermography, a nondestructive testing method for composite materials, is improved in this study with a Sparse Principal Component Thermography (SPCT) method (wu2018sparse, ). Building on the advantages of Principal Component Thermography (PCT), SPCT introduces a penalization term, enhancing flexibility and interpretability due to its structural sparsity. The effectiveness of SPCT is proven through experiments on a carbon fibre-reinforced plastic specimen, demonstrating superior subsurface defect visualization compared to PCT. Machine learning methods are increasingly significant in the field of nondestructive testing, particularly for the quality assessment of polymer composites. Among these, a Generative Adversarial Network (GAN), a form of deep learning, is introduced to thermography as an image augmentation tool, enhancing defect detection capabilities (liu2020generative, ). This method uses data augmentation to produce more diverse and informative images, improving the defect detection capabilities of thermographic analysis. The effectiveness of GPCT is demonstrated through successful experiments on a carbon fibre-reinforced polymer specimen, showcasing its potential in nondestructive testing.
3.1.4. Self-organizing Maps
@@
Self-organizing Maps (SOMs) are a type of unsupervised neural network used for data visualization and pattern recognition, which can be effectively applied to material defect detection (shakhovska2020improved, ). An automated visual inspection method for detecting and classifying textile defects, using the autocorrelation function for feature extraction and a Self-Organizing Feature Map (SOFM) (tolba1997self, ). This approach is immune to illumination changes and noise, thanks to the noise rejection property of the autocorrelation function. Tested on real textile samples, the SOFM effectively clusters the feature vectors, showcasing the advantages of unsupervised learning systems in defect classification. For textured pavement images, Mathavan et al. (mathavan2015use, ) study a method using a self-organizing map (SOM) for detecting road cracks in textured pavement images. It involves analyzing images segmented into cells’ texture and colour properties and using a Kohonen map to identify cracks. The technique achieves 77% precision and 73% recall, showing effectiveness even in uncontrolled environments and outperforming an advanced algorithm in reducing false positives in crack detection.
3.1.5. Variational Autoencoders
@@
Variational Autoencoders (VAEs) are a type of deep learning model particularly well-suited for anomaly and defect detection in materials. Using unsupervised learning, Tsai et al. (tsai2021autoencoder, ) evaluate an improved Convolutional Autoencoder (CAE) model for automated defect detection in manufacturing. Unlike traditional CAEs that minimize reconstruction errors, this enhanced CAE incorporates a regularization to tighten the feature distribution of defect-free samples. This approach makes defect samples stand out more distinctly during evaluation. Tested on various material surfaces, including textured and patterned ones, the modified CAE demonstrates superior performance in defect detection compared to conventional CAEs. In addition, a variational autoencoder-enhanced deep learning model (VAEDLM) for efficiently classifying imbalanced wafer defects in semiconductor foundries (wang2021variational, ). The model effectively generates and recognizes wafer defect patterns by utilizing variational autoencoders and a refined convolutional neural network. Tested on the WM-811K dataset, the VAEDLM achieves exceptional performance metrics including 99.19% accuracy and 99.96% AUC, outperforming current methods. Additionally, the model’s interpretability is enhanced through saliency maps and t-SNE, addressing a common limitation in deep learning applications.
3.1.6. Deep Belief Network
@@
A Deep Belief Network (DBN) is a type of deep neural network with multiple layers of stochastic, latent variables. It’s particularly useful for feature extraction and classification tasks, making it suitable for material defect detection. Cang et al. (cang2017microstructure, ) focus on enhancing Integrated Computational Materials Engineering (ICME) by introducing a Convolutional Deep Belief Network (CDBN) for feature learning. This approach automates the transformation of microstructures into lower-dimensional representations, achieving a significant reduction in dimensionality. The model is applied to diverse material systems, including alloys and sandstone, successfully reconstructing materials closely resembling the originals in terms of key properties. This method outperforms traditional techniques that rely on the Markovian assumption, showcasing its potential in advanced material design and reconstruction. Furthermore, Liu et al. (liu2019real, ) introduce a real-time quality monitoring and diagnosis scheme for manufacturing processes using a DBN. It addresses the challenge of handling complex, high-dimensional process data by converting them into quality spectra, which are then analyzed using a DBN. This model is developed through offline learning and applied for online process monitoring and diagnosis. Its effectiveness is demonstrated through simulation experiments and a practical example in injection moulding, showing superior performance compared to traditional methods.
@@ 3.1.7. Unsupervised Technique Summary
The application of unsupervised learning techniques in material defect detection exhibits diversity and versatility. Table 2 summarizes the application scenarios, advantages, and limitations of various unsupervised techniques. As shown in Table 2, Histogram-based techniques are applicable for detecting specific types of defects, such as nodular defects on wood surfaces, but may not be suitable for other types of defects. Clustering algorithms are capable of detecting material defects by grouping similar data points, offering wide applicability for identifying anomalies or defects, albeit potentially requiring a large amount of unlabeled data to enhance accuracy. PCA simplifies the complexity of high-dimensional data, yet may be insufficient for handling more complex defect types. SOMs are utilized for data visualization and pattern recognition, effectively classifying feature vectors, but may yield suboptimal results outside controlled environments. VAEs highlight defect samples and are suitable for detecting anomalies and defects in materials, yet may necessitate a significant amount of training data. DBNs automatically transform microscopic structures and are employed for feature extraction and classification tasks, though they may require substantial training and computational resources. Considering task requirements, data characteristics, and computational resources, selecting the appropriate technique is crucial for improving the accuracy and efficiency of material defect detection.
Table 2. Unsupervised Technique Comparison in Material Defect Detection
Unsupervised Technique Application Scenarios Advantages Limitations
@@
specific types of defects, such as
nodular defects on wood surfaces
effective identification of specific
types of defects
not applicable to
other types of defects
Clustering algorithms multiple similar detects identify generic anomalies or defects
require a large amount of
@@
Principal Component Analysis
simplify the complexity of
high-dimensional
detect feature data
retain defect trends and patterns
not sufficient to handle more
complex defect types
Self-organizing Maps
defect data visualization
and pattern recognition
effectively classify detect
feature vectors and
insensitive to the distribution
and features of defect data
highly sensitive to parameter
selection and high
computational complexity
Variational Autoencoders
Scarcity of defect samples,
anomaly recognition
in defect detection.
enrich training samples of defects
and identify anomalies
in defect detection.
require a large amount
of training data
Deep Belief Network
defect feature extraction
@@
identify anomalies
in defect detection.
require significant training
and computational resources
@@ 3.2. Supervised Learning Techniques in Material Defect Detection
This section summarizes the supervised learning techniques relevant to material defect detection. Fig. 9 outlines a general workflow for this process, divided into five main steps: data collection, data labelling, data preprocessing, model selection and training, and model testing. Initially, a sensor captures data to form the original dataset. This dataset is then labelled with defect categories in the data labelling step. Next, the data undergoes preprocessing, which includes data cleaning, augmentation, and normalization, resulting in a processed dataset. In the model selection and training step, a designed learning model is trained using forward and backward propagation methods on the processed dataset. Finally, the well-trained model is tested with test samples to detect and classify material defects.
@@
3.2.1. Regression Algorithms
Regression algorithms are a class of statistical methods used in machine learning for predicting a continuous outcome. They can be effectively applied to material defect detection, particularly in predicting the extent, severity, or type of defects based on various input features. To detect the defection efficiently, a two-stage neural algorithm for defect detection and depth characterization using active thermography is proposed (dudzik2015two, ). The first stage involves a classification neural network for detecting defects, followed by a regression neural network to estimate their depth. Results from experimental investigations and simulations are presented, along with a sensitivity analysis that examines the effects of emissivity and ambient temperature errors on depth estimation. The algorithm’s performance is validated using a test sample of low thermal diffusivity material. Various regression algorithms have been utilized in detecting defects during the Laser Beam Additive Manufacturing (LBAM) process. In their work, Mahmoudi et al. developed a system for detecting anomalies specifically in the Laser Powder Bed Fusion (LPBF) printing process (mahmoudi2019layerwise, ).
@@ 3.2.2. Neural Network Algorithms
For a long time, the practical use of Thermographic Non-Destructive Evaluation (TNDE) has faced challenges due to the intricate and nonlinear nature of related inversion problems, such as accurately determining the depth of defects. However, artificial neural networks have recently shown promising results in detecting and quantitatively assessing defects in TNDE. Darabi et al. (darabi2002neural, ) introduce a three-dimensional thermal model for non-homogeneous materials, like carbon fibre-reinforced plastic (CFRP). The outcomes of this model are then compared with analytical solutions derived from Duhamel’s theorem. Subsequently, the paper describes the development of two back-propagation neural networks, one for detecting defects and the other for estimating their depth. Lastly, it presents and examines both simulated and experimental results. Furthermore, to improve the detection performance, an advanced convolutional neural network-based method for accurately detecting surface defects in workpieces has been proposed (xing2021convolutional, ). It introduces a specialized model (SCN) for feature extraction and uses an optimized IOU (XIoU) for loss function definition. The method demonstrates superior performance over other models in both classification and detection tasks, with high accuracy and real-time detection capabilities, making it highly effective for automatic surface defect detection in workpieces.
3.2.3. Ensemble Learning Algorithms

### v3

@@ 3.1.1. Feature-Based Techniques
PCA is a statistical technique used in MDD to simplify the complexity of high-dimensional data while retaining trends and patterns (cao2024crack, ). Active thermography, a nondestructive testing method for composite materials, is improved in this study with a Sparse Principal Component Thermography (SPCT) method (wu2018sparse, ). Building on the advantages of Principal Component Thermography (PCT), SPCT introduces a penalization term, enhancing flexibility and interpretability due to its structural sparsity. The effectiveness of SPCT is proven through experiments on a carbon fibre-reinforced plastic specimen, demonstrating superior subsurface defect visualization compared to PCT. ML methods are increasingly significant in the field of nondestructive testing, particularly for the quality assessment of polymer composites. Among these, a Generative Adversarial Network (GAN), a form of DL, is introduced to thermography as an image augmentation tool, enhancing defect detection capabilities (liu2020generative, ). This method, termed GPCT, uses data augmentation to produce more diverse and informative images, improving the defect detection capabilities of thermographic analysis. The effectiveness of GPCT is demonstrated through successful experiments on a carbon fibre-reinforced polymer specimen, showcasing its potential in nondestructive testing.
Self-Organizing Maps (SOMs)-Based
@@
SOMs are a type of unsupervised neural network used for data visualization and pattern recognition, which can be effectively applied to MDD (shakhovska2020improved, ). An automated visual inspection method for detecting and classifying textile defects, using the autocorrelation function for feature extraction and a Self-Organizing Feature Map (SOFM) (tolba1997self, ). This approach is immune to illumination changes and noise, thanks to the noise rejection property of the autocorrelation function. Tested on real textile samples, the SOFM effectively clusters the feature vectors, showcasing the advantages of UL systems in defect classification. For textured pavement images, Mathavan et al. (mathavan2015use, ) study a method using a SOM for detecting road cracks in textured pavement images. It involves analyzing images segmented into cells’ texture and colour properties and using a Kohonen map to identify cracks. The technique achieves 77% precision and 73% recall, showing effectiveness even in uncontrolled environments and outperforming an advanced algorithm in reducing false positives in crack detection.
Figure 4. Illustration of the general working process for conducting the material defect detection with UL techniques.
3.1.2. Clustering-Based Techniques
Clustering techniques group data into subsets based on feature similarities, making them effective for anomaly detection in material defects. These techniques can be broadly categorized into centroid-based, hierarchical, and hybrid clustering methods, each with unique strengths in MDD.
Centroid-Based Clustering (e.g., K-Means Clustering)-Based
K-means clustering is widely applied in defect segmentation and clustering due to its simplicity and efficiency. For example, (cohn2021unsupervised, ) achieved 99.4% accuracy in classifying steel surface defects using features refined by PCA and extracted through a pre-trained VGG16 model. PCA reduced dimensionality, and K-means classified defects without requiring labeled data, offering valuable insights into unsupervised learning for materials science. Similarly, Thomas et al. (leung2001representing, ) proposed a centroid-based clustering method for material recognition using 3D textons, compact surface patches encoding geometric and photometric properties. These textons, clustered into a vocabulary of representative patterns, enabled effective material recognition and appearance prediction under varied conditions.
Hierarchical Clustering-Based
Hierarchical clustering constructs a hierarchy of clusters, enabling multi-scale defect detection. Jorge et al. (igual2020hierarchical, ) present a hierarchical clustering approach using impact-echo techniques for defect detection in materials. Feature vectors, indicating wave propagation changes caused by defects, are analyzed as mixtures of Gaussians. Clusters are merged using the Kullback-Leibler divergence. Applied to aluminium alloy specimens, the method successfully classifies them into categories like no defect, one hole, one crack, and multiple defects, effectively distinguishing between defective and non-defective materials. Brzakovic et al. (brzakovic1990approach, ) introduce an expert system for identifying and categorizing defects in materials with complex textures using digitized images. It involves a two-step image segmentation process and a material-specific defect classification approach. A hierarchical classification scheme incorporating shape, size, and texture descriptors is used for wood. The system, tested on parquet samples, has successfully detected and classified various defects.
Hybrid Clustering (e.g., Fuzzy and Spectral Clustering)-Based
Hybrid clustering combines different clustering methods to handle more complex defect patterns. For example, in high-tech materials such as semiconductor manufacturing, a spatial defect diagnosis system has been proposed to estimate defect clusters and their separation, including both convex and non-convex types. The approach (wang2008recognition, ) uses a spatial filter to clean wafer maps and a hybrid of entropy fuzzy c-means and spectral clustering to isolate defect patterns. A decision tree based on cluster features helps identify defect types. Tested with data from a DRAM company in Taiwan, the method effectively extracts and classifies various defect patterns, showing potential for application in other industries like liquid crystal and plasma display manufacturing.
3.1.3. Deep Learning (DL)-Based Techniques
DL-based techniques in UL are designed to automatically learn high-level feature representations from raw data, capturing intricate patterns and structures that traditional methods might overlook. These models, such as Variational Autoencoders (VAEs) and Deep Belief Networks (DBNs), excel in identifying complex and non-linear relationships within the data, making them particularly effective for anomaly detection and defect characterization in MDD workflows.
Variational Autoencoders (VAEs)-Based
@@ 3.1.3. Deep Learning (DL)-Based Techniques
VAEs are a type of DL model particularly well-suited for anomaly and defect detection in materials. Using UL, Tsai et al. (tsai2021autoencoder, ) evaluate an improved CAE model for automated defect detection in manufacturing. Unlike traditional CAEs that minimize reconstruction errors, this enhanced CAE incorporates a regularization to tighten the feature distribution of defect-free samples. This approach makes defect samples stand out more distinctly during evaluation. Tested on various material surfaces, including textured and patterned ones, the modified CAE demonstrates superior performance in defect detection compared to conventional CAEs. In addition, a VAE-enhanced DL model (VAEDLM) for efficiently classifying imbalanced wafer defects in semiconductor foundries (wang2021variational, ). The model effectively generates and recognizes wafer defect patterns by utilizing VAEs and a refined CNN. Tested on the WM-811K dataset, the VAEDLM achieves exceptional performance metrics, including 99.19% accuracy and 99.96% AUC, outperforming current methods. Additionally, the model’s interpretability is enhanced through saliency maps and t-SNE, addressing a common limitation in DL applications.
Deep Belief Networks (DBNs)-Based
@@
A DBN is a type of deep neural network with multiple layers of stochastic, latent variables. It’s particularly useful for feature extraction and classification tasks, making it suitable for MDD. Cang et al. (cang2017microstructure, ) focus on enhancing Integrated Computational Materials Engineering (ICME) by introducing a Convolutional DBN (CDBN) for feature learning. This approach automates the transformation of microstructures into lower-dimensional representations, achieving a significant reduction in dimensionality. The model is applied to diverse material systems, including alloys and sandstone, successfully reconstructing materials closely resembling the originals in terms of key properties. This method outperforms traditional techniques that rely on the Markovian assumption, showcasing its potential in advanced material design and reconstruction. Furthermore, Liu et al. (liu2019real, ) introduce a real-time quality monitoring and diagnosis scheme for manufacturing processes using a DBN. It addresses the challenge of handling complex, high-dimensional process data by converting it into quality spectra, which are then analyzed using a DBN. This model is developed through offline learning and applied to online process monitoring and diagnosis. Its effectiveness is demonstrated through simulation experiments, showing superior performance compared to traditional methods.
3.1.4. Self-Supervised Learning Techniques
Self-supervised learning (SSL) has emerged as a transformative approach within the realm of unsupervised techniques for MDD. SSL leverages unlabeled data by employing pretext tasks to generate pseudo-labels, enabling models to learn meaningful representations without manual annotations. This characteristic is particularly beneficial in scenarios where labeled defect data is scarce or expensive to obtain. For instance, Zhang et al. (zhang2024maegan, ) proposed a Masked Autoencoder (MAE)-GAN model for cigarette defect detection, achieving robust performance in appearance defect identification. Similarly, Era et al. (era2024thermal, ) employed SSL with Vision Transformers for melt pool characterization in directed energy deposition, enhancing thermal imaging-based defect detection. Anomaly detection in photovoltaic arrays using a dual-source SSL discriminator further demonstrates the method’s versatility in handling noisy environments (yao2024anomaly, ). Additionally, Li et al. (li2024steel, ) introduced a single-class SSL framework for surface defect detection in steel bars, effectively addressing imbalanced data distributions. Beyond these applications, Hu et al. (hu2024segmentation, ) utilized SSL combined with multi-task learning to enhance defect segmentation accuracy, while Helwing et al. (helwing2024damage, ) applied cycle-consistent GANs for damage evolution analysis in fiber-reinforced polymers. These advancements underscore SSL’s ability to address complex defect scenarios by autonomously extracting and refining features, making it an indispensable tool for advancing MDD.
@@ 3.1.5. Unsupervised Technique Summary
Table 4. Comparison of unsupervised learning techniques for material defect detection
Technique Category Application Scenarios Advantages Limitations
Feature-Based Techniques
Specific defect types,
data simplification
Simplifies high-dimensional
data; highlights key
features
Limited to specific
feature extraction tasks
@@
Detecting specific
defect types
Effective for targeted
defect identification
Not generalizable to
all defect types
- Principal Component Analysis (PCA)
Preprocessing high-
dimensional data
Simplifies data;
retains key patterns
Not suitable for
complex defect types
- Self-Organizing Maps (SOMs)
Pattern recognition
and data visualization
Effective for clustering
feature vectors
Sensitive to parameters;
computationally intensive
Clustering-Based Techniques
Anomaly detection
and defect grouping
Groups similar data
points effectively
Requires significant
@@
- Centroid-Based Clustering (K-Means)
Defect segmentation
and clustering
Simple and efficient for
clustering similar defects
Less effective for
non-linear relationships
- Hierarchical Clustering
Multi-scale defect
detection
Handles multi-scale
defect patterns
High computational
complexity
- Hybrid Clustering
Complex defect
pattern detection
Combines strengths of
multiple clustering methods
Computationally
expensive
Deep Learning-Based Techniques
Anomaly detection
and defect characterization
Automatically learns
high-level features
Requires extensive
training and data
- Variational Autoencoders (VAEs)
Anomaly detection
in scarce data
Highlights anomalies;
generates diverse samples
Depends heavily on
training data
- Deep Belief Networks (DBNs)
Feature extraction
@@
Captures complex
defect features
Computationally
intensive
Self-Supervised Learning (SSL)
Learning meaningful
representations from
unlabeled data
Reduces reliance on
labeled data; enhances
feature extraction
Requires complex pretext
tasks and substantial
computational resources
The application of UL techniques in MDD exhibits diversity and versatility. Table 4 summarizes the application scenarios, advantages, and limitations of various unsupervised techniques. As shown in Table 4, feature-based techniques, including histogram-based methods, PCA, and SOMs, are commonly used for extracting key characteristics from raw data, enabling simplified representations for further analysis. Histogram-based techniques effectively detect specific types of defects, such as nodular defects on wood surfaces, but may not generalize to other defect types. PCA is valuable for reducing the complexity of high-dimensional data while retaining critical patterns, yet it is insufficient for addressing more complex defect scenarios. SOMs are useful for data visualization and clustering, showcasing their strength in pattern recognition, though their performance can be sensitive to parameter settings and environmental variability. Clustering-based techniques, such as K-Means, hierarchical clustering, and hybrid clustering, are highly effective for identifying defect patterns and grouping similar data points. These techniques offer broad applicability but often require significant amounts of unlabeled data for optimal performance. DL-based techniques, including VAEs and DBNs, excel in learning high-level representations and detecting anomalies or complex defect features. VAEs are particularly suitable for generating diverse samples and highlighting anomalies, making them effective in cases with scarce defect data, but they require extensive training data. DBNs are powerful tools for feature extraction and defect classification, but are computationally intensive and demand significant training resources. Additionally, SSL techniques have gained traction for leveraging unlabeled data through pretext tasks, enabling models to learn meaningful representations and improving defect detection in scenarios with limited labeled samples. Considering task requirements, data characteristics, and computational constraints, selecting the appropriate technique is crucial for improving the accuracy and efficiency of MDD.
@@ 3.2. Supervised Learning Techniques in Material Defect Detection
This section summarizes SL techniques relevant to MDD. Fig. 5 outlines a general workflow for this process, divided into five main steps: data collection, data labeling, data preprocessing, model selection and training, and model testing. Initially, a sensor captures data to form the original dataset. This dataset is then labeled with defect categories in the data labeling step. Next, the data undergoes preprocessing, which includes data cleaning, augmentation, and normalization, resulting in a processed dataset. In the model selection and training step, a designed learning model is trained using forward and backward propagation methods on the processed dataset. Finally, the well-trained model is tested with test samples to detect and classify material defects.
@@
3.2.1. Regression-Based Techniques
Regression-based techniques are statistical and ML methods used for predicting continuous outcomes, such as defect depth, severity, or extent. These methods are commonly applied in scenarios requiring quantitative defect characterization. To detect the defect efficiently, a two-stage neural algorithm for defect detection and depth characterization using active thermography is proposed (dudzik2015two, ). The first stage involves a classification neural network for detecting defects, followed by a regression neural network to estimate their depth. Results from experimental investigations and simulations are presented, along with a sensitivity analysis that examines the effects of emissivity and ambient temperature errors on depth estimation. The algorithm’s performance is validated using a test sample of low thermal diffusivity material. Various regression algorithms have been utilized in detecting defects during the Laser Beam Additive Manufacturing (LBAM) process. In their work, Mahmoudi et al. developed a system for detecting anomalies specifically in the Laser Powder Bed Fusion (LPBF) printing process (mahmoudi2019layerwise, ).
@@ 3.2.2. Neural Network Algorithms
Neural network algorithms play a critical role in MDD, leveraging diverse structures to address complex, nonlinear problems. These architectures are broadly categorized based on their design and application focus:
Convolutional Neural Networks (CNNs)
CNNs are specialized for processing structured grid data, such as images, making them ideal for surface defect detection. Xing et al. (xing2021convolutional, ) proposed an advanced CNN-based method for accurately detecting surface defects in workpieces. This model integrates a specialized feature extraction module and uses an optimized IOU-based loss function to enhance performance. The method demonstrated superior results over other models in both classification and detection tasks, achieving high accuracy and real-time detection capabilities. CNNs have thus become indispensable tools in industrial defect detection applications.
Attention Mechanisms (Modules)-Based
Attention mechanisms, as modules integrated within neural network frameworks, enable models to focus on specific parts of the input data relevant to the defect detection task. These mechanisms significantly enhance the accuracy and efficiency of defect detection, particularly in challenging environments. For example, Fang et al. (fang2022tactile, ) introduced a CNN-based method incorporating attention mechanisms for detecting fabric defects, using a vision-based tactile sensor. This approach improved detection robustness against ambient noise and complex dyeing patterns. Similarly, Chen et al. (chen2021defect, ) utilized a self-attention mechanism for defect detection on aluminum profile surfaces, achieving superior performance in noisy conditions by refining feature extraction and employing a residual learning strategy.
Transformer Models
Transformers, such as Vision Transformers (ViT), have emerged as powerful tools for defect detection by capturing global context and relationships in data. Liu et al. (liu2024multi, ) proposed a Multi-feature Vision Transformer for automatic defect detection in composites, which integrates thermography with advanced feature representation. Wang et al. (wang2023defect, ) developed Defect Transformer, a hybrid architecture combining CNN properties with transformer structures to enhance pixel-level defect detection precision in industrial materials like steel. These studies highlight transformers’ ability to process high-resolution images and complex material characteristics.
Graph Neural Networks (GNNs)
GNNs are increasingly used for analyzing structured data in material science. Pang and Tan (pang2024steel, ) developed a steel surface defect detection model based on GNNs, leveraging graph representations to improve detection precision and efficiency. Thomas et al. (thomas2023materials, ) applied GNNs for material fatigue prediction using microstructure representations, demonstrating their capability in capturing complex relationships within material microstructures. Additionally, Wang et al. (wang2022graph, ) proposed a graph-guided CNN (G-GCNN) for surface defect recognition, which effectively combines graph-based representations and convolutional features to identify complex defect patterns. These findings showcase GNNs’ versatility in modeling structured data and their effectiveness in industrial defect detection tasks.
Recurrent Neural Networks (RNNs)
RNNs, particularly their variants such as Long Short-Term Memory (LSTM) networks, are effective in time-series defect detection, such as monitoring equipment performance or dynamic surface changes. These architectures capture temporal dependencies and trends critical for predictive maintenance. For example, Hu et al. (hu2019lstm, ) applied LSTM-RNN to classify defects in honeycomb structures using infrared thermography, demonstrating high accuracy and efficiency in handling complex thermal data. Similarly, Wang et al. (wang2020defect, ) proposed an LSTM-RNN-based method for determining defect depth in laser infrared thermography, achieving precise defect depth predictions.
Hybrid Neural Networks
Hybrid neural networks integrate features from multiple architectures, such as CNNs, transformers, and attention mechanisms, to address complex defect detection tasks, especially in multimodal industrial scenarios. These networks leverage complementary strengths of different models to enhance detection precision and robustness. For instance, Zhao et al. (zhao2024deep, ) developed a hybrid approach combining deep CNNs and multi-model fusion for optical inspection of hybrid bonded wafer defects. This method significantly improved detection reliability and accuracy for defects affecting hybrid bonding yields. Similarly, Zhou et al. (zhou2023defect, ) introduced a hybrid attention network that integrates CNNs and auto-encoders (AE) for steel defect detection. Their model localized and classified defect regions with high precision, demonstrating robustness in challenging industrial conditions. These advancements underscore the potential of hybrid neural networks in achieving superior performance in diverse MDD applications.
3.2.3. Ensemble Learning-Based Techniques